OpenStudy (anonymous):

Find x and y. Remember that a + bi = c + di exactly when a = c and b = d. 8 = (4+8i)(x+yi)

OpenStudy (lalaly):

\[x+yi = \frac{8}{4+8i}\]\[x+yi=\frac{2}{1+2i}\]\[x+yi=\frac{2}{1+2i} \times \frac{1-2i}{1-2i}\]\[x+yi=\frac{2-4i}{5}\]\[x+yi=\frac{2}{4}-\frac{4}{5}i\]

OpenStudy (lalaly):

one mistake last step \[x+yi=\frac{2}{5} - \frac{4}{5}i\]

OpenStudy (lalaly):

\[x=\frac{2}{5}\]\[y=\frac{-4}{5}\]

OpenStudy (anonymous):

wolfram says .... agree ! http://www.wolframalpha.com/input/?i=8+%3D+%284%2B8i%29*X

OpenStudy (lalaly):

yaaay x)